OpenStudy (anonymous):

Help please! 1) A comet that passes through the solar system just once has a path that is modeled by a conic section. Astronomers discover a comet whose path is modeled by the following equation, with the Sun at one focus. What type of conic section is the comet's path? x^2/2,500-y^2/37,500=1 2) What are the vertices of the comet's path? (Hint: Examine the general formula for the path and identify the numbers that determine the vertices.) 3) At its closest point to the Sun, this comet is ____ million miles from the Sun. (Enter only the number.) Hint: Subtract the vertex from the focus Thank you!

OpenStudy (mertsj):

1. Hint: The negative sign gives it away.

OpenStudy (anonymous):

Negative sign? @Mertsj

OpenStudy (mertsj):

Yes. do you see the minus sign between the terms on the left side?

OpenStudy (anonymous):

On number 1? Yes @Mertsj

OpenStudy (mertsj):

It identifies the conic section for you if you look at the general equations.

OpenStudy (mertsj):

Which conic section is it?

OpenStudy (anonymous):

Um... ellipse?

OpenStudy (mertsj):

http://math2.org/math/algebra/conics.htm

OpenStudy (mertsj):

Scroll to the bottom of that page and look for the one that looks like your equation.

OpenStudy (anonymous):

Hyperbola? @Mertsj

OpenStudy (mertsj):

Good for you!!

OpenStudy (anonymous):

Wooo hooo! :) @Mertsj Can you help with the other two please?

OpenStudy (mertsj):

Before you find the vertices, you need to identify the center. Can you do that?

OpenStudy (anonymous):

Um... I don't think so. I have a hard time with math.

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}-\frac{(x-k)^2}{b^2}=1\]

OpenStudy (mertsj):

In that equation (h,k) is the center. Can you match your equation up with that one and identify h and k?

OpenStudy (anonymous):

Would they just both be 1?

OpenStudy (mertsj):

I think the reason they aren't there is because they are both 0. So the center is (0,0)

OpenStudy (anonymous):

Oh okay

OpenStudy (mertsj):

And now you need to identify a^2

OpenStudy (anonymous):

a^2 is 2,500

OpenStudy (mertsj):

a^2 is the number under x^2. what is it?

OpenStudy (mertsj):

And so yes it is 2500 so what is a?

OpenStudy (anonymous):

50

OpenStudy (mertsj):

Yes. Now a is important because the vertices are a units to the right and a units to the left of the center. So what are the vertices?

OpenStudy (anonymous):

(50, 0) and (-50, 0)

OpenStudy (mertsj):

Good.

OpenStudy (anonymous):

Okay. :)

OpenStudy (mertsj):

Now we need to identify b.

OpenStudy (mertsj):

What is it?

OpenStudy (anonymous):

Um... i got 193.64

OpenStudy (mertsj):

me too.

OpenStudy (mertsj):

Now find c by using the equation: a^2+b^2=c^2

OpenStudy (mertsj):

We need c so we can find the foci.

OpenStudy (anonymous):

I got 40000

OpenStudy (mertsj):

That would be c^2. Right?

OpenStudy (anonymous):

Yeah

OpenStudy (mertsj):

So c is?

OpenStudy (anonymous):

200

OpenStudy (mertsj):

c is important because the foci are c units to the right and c units to the left of the center.

OpenStudy (mertsj):

And now you can apply the hint given in 3.

OpenStudy (anonymous):

Wait... hold on... I'm confused

OpenStudy (mertsj):

The hint says to subtract the vertex from the focus. What is 200-50?

OpenStudy (anonymous):

Oh. 150

OpenStudy (mertsj):

So the closest the comet comes to the sun is 150 million miles.

OpenStudy (anonymous):

Okay. Cool! If I could hug you right now I would! You are so awesome! Thanks!!!!

OpenStudy (mertsj):

You are welcome.